OpenStudy (august899):

The point (2,n) lies on the circle whose equation is (x-3)^2+(y+2)^2=26.Find the value of n.

hartnn (hartnn):

you can plug in x= 2, and y =n in the circle's equation. and solve for n :)

OpenStudy (seratul):

If the x is 2, you can easily plug it in the equation and solve for it.

OpenStudy (seratul):

Oh, yea what hartnn said :P

hartnn (hartnn):

note: you'll get 2 values of n

OpenStudy (august899):

So it would be (2-3)^2+(y+2)^2=26?

hartnn (hartnn):

yes

OpenStudy (august899):

Okay,then do I just do the 2-3^2 and get -1 then do -1+(y+2)^2=26?

hartnn (hartnn):

brackets are important :) \((2-3) = -1 \\ (2-3)^2 = (-1)^2 = +1\)

hartnn (hartnn):

\((-1)^2 = (-1)\times(-1) = 1\) Also, without brackets, 2-3^2 = 2-9 = -7 which is totally incorrect :P

OpenStudy (dinamix):

\[+1+n^2+4n+4=26\] \[n^2+4n=21\] \[n^2+4n-21 =0\]

hartnn (hartnn):

doing square root on both sides would be easier :)

OpenStudy (dinamix):

@August899 use delta to find n that's simple

OpenStudy (august899):

So I do -1+(y+2)^2=26,What do I do afterwards?

hartnn (hartnn):

its not -1 \((2-3)^2 + (n+2)^2 = 26 \\ (-1)^2 + (n+2)^2 = 26 \\1 + (n+2)^2 = 26\) you get this?

OpenStudy (august899):

Oh okay. Yes I get it now.

hartnn (hartnn):

now subtract 1 from both sides

OpenStudy (august899):

Okay.So now it is (n+2)^2=25

hartnn (hartnn):

yes take square root on both sides

OpenStudy (august899):

I square root the (n+2)^2 correct?

OpenStudy (august899):

and that equals 5 because the square root of 25 is 5

hartnn (hartnn):

with your method, dinamix, b^2 -4ac = 16- 4(1)(-21) = 16+4*21 = ... yes August.

OpenStudy (dinamix):

yup sorry its +17

hartnn (hartnn):

@August899 \(a^2 = b^2 \implies a = \pm b\)

hartnn (hartnn):

\((n+2)^2 =5^2 \implies n+2 = \pm 5\)

OpenStudy (august899):

How do I do the square root of (n+2)^2? Do I do 2^2 which 4 and do the square root which is 2 and do 2=25 and divide by 2?

hartnn (hartnn):

\(\Large \sqrt{(anything)^2 } = anything \\ \Large \sqrt{(n+2)^2} = n+2\)

OpenStudy (august899):

Ohh okay.Then I subtract 2 from both sides?

OpenStudy (zarkon):

\[\sqrt{(n+2)^2}=|n+2|\]

hartnn (hartnn):

yes yes

OpenStudy (august899):

Okay I got n=3

OpenStudy (august899):

Thank you sooo much for the help!

hartnn (hartnn):

\(\color{blue}{\text{Originally Posted by}}\) @Zarkon \[\sqrt{(n+2)^2}=|n+2|\] \(\color{blue}{\text{End of Quote}}\) \(|n+2| = 5 \\n+2 = \pm 5 \)

hartnn (hartnn):

you will have one more value of n

hartnn (hartnn):

n +2 = 5 ... n = 3 n+2 = -5 ... n = ...?

OpenStudy (august899):

7?

hartnn (hartnn):

n +2 = -5 subtract 2 from both sides

OpenStudy (august899):

-7

hartnn (hartnn):

\(\checkmark \)

hartnn (hartnn):

so the 2 values of n are 3, -7 :)

OpenStudy (august899):

Thank you! I understand it now.

hartnn (hartnn):

welcome ^_^